# Accurate Hyperfine Tensors for Solid State Quantum Applications: Case of the NV Center in Diamond

### v2

@@ Introduction
Hyperfine coupling tensors can be calculated using different first principles methods, such as density functional theory (DFT) and wave functions-based methods. The accuracy of the computed hyperfine tensors for close nuclear spins is remarkable. For example, considering paramagnetic point defects in semiconductors a mean absolute relative error of 4.7% has been reported.Szász et al. (2013) Since the hyperfine structure of the point defect’s spin sublevels is unique like a bar code, one can compare the measured and computed hyperfine tensors to unambiguously identify paramagnetic defects in semiconductors, see for instance Refs. Son et al. ; Isoya et al. ; Ivády et al. (2017).
@@
As demonstrated recently by a supercell-size-scaling test in Ref. Ivády et al. (2018), the remarkable accuracy of the computed hyperfine parameters is only limited for the closest nuclear spin in 1-5 Å distances from the defect. The relative error sharply increases for nuclear spins located at large distances. As discussed in Ref. Ivády et al. (2018), this is presumably due to the periodic boundary condition and related finite size effects. Using finite cluster models, such as the CH cluster and the CH cluster used in Refs. Nizovtsev et al. (2014, ) to calculate the hyperfine coupling tensors at 1.5-8 Å distances from the defect, errors related to the periodic boundary condition can be eliminated. Another limitation of first principles calculations is the model size that maximizes the number of lattice sites that can be considered.
@@
In this work, we first demonstrate the inaccuracy of the numerical hyperfine parameters obtained with the industry standard VASP codeKresse and Hafner (1994); Kresse and Furthmüller (1996). To resolve the underlying methodological issues we introduce a real-space integration method and the use of a large support lattice for considering nuclear spins outside the boundaries of the supercell. To benchmark our method, we carry out large-scale calculations for the NV center in diamond using different exchange-correlation functionals and compare the numerical results with available experimental hyperfine data sets. From the comparison, we conclude that the HSE06Heyd et al. (2003, 2006) functional with 0.2 mixing parameter performs best for the NV center in diamond resulting in a mean absolute percentage error of 1.7% for nuclear spins 6-30 Å distances from the NV center. This is a significant improvement compared to previous theoretical predictions obtained by using VASP. We show that the residual errors are likely related to the inaccurate calculation of the Fermi contact term. High accuracy hyperfine tensors of lattice sites as well as volumetric hyperfine data with Å spatial resolution are published together with this Letter at Ref. 111Theoretical hyperfine tensors for all lattice sites within 30 Å distance from the NV center are available at https://ivadygroup.elte.hu/hyperfine and ready to be used for modeling NV center quantum nodes and positioning nuclear spins around the NV center in diamond in nano-NMR measurements.
@@ I Results and Discussion
where is the hyperfine tensor and and are the electron spin and the nuclear spin operator vectors with and quantum numbers for the NV center and adjacent C nuclear spins of the diamond lattice, respectively. The two dominant contributions to the hyperfine tensor are the Fermi contact interaction, , and the magnetic dipole-dipole coupling of the electron and nuclear spins, .
@@
First, we use this method as implemented in VASP to calculate hyperfine tensors of all sites in a 512-atom and a 1728-atom supercell of diamond including a single NV center in the middle of the supercell, see Fig. 1. For the calculations, we use HSE06 exchange-correlation functional, 500 eV cutoff energy of the plane-wave basis set, -point sampling of the Brillouin zone, and high convergence criteria. The structure of the defect is optimized as far as the largest force is smaller than 10 eV/Å. To compare our results with the experimental values, see Fig. 2(a), we either compute the hyperfine splitting as Dréau et al. (2012), where are the elements of the hyperfine tensor, which is compared with experimental values in data set I. taken from Ref. Dréau et al. (2012), or compare the hyperfine tensor element with experimental values obtained from high precision measurements in Ref. Taminiau et al. (2012) (data set II.) and Ref. van de Stolpe et al. (2023) (data set III.). Note that the latter data set is a refined and extended version of the data published in Ref. Abobeih et al. (2019). From the reported 50 nuclear spins, we consider only those for which both the position and the hyperfine parameter are reported with high accuracy, see Fig. 2(a). When comparing the computed values with data sets II.-III., we found a consistent sign difference between theory and experiment. The sign of our hyperfine values agrees with other theoretical calculations, i.e. VASP’s values as well as the values reported in Refs. Nizovtsev et al. (2014, ). Therefore, we anticipate that the discrepancy originates from the convention used in the experiments. Hereinafter, we omit this sign difference.
@@
The hyperfine splitting values computed with our method are significantly improved compared to the currently available implementationSzász et al. (2013). The relative error of the theoretical values reduce from to , see Fig. 2(b) and Fig. 3. For different data sets, we obtain different mean absolute percentage errors (MAPE). For data set I. we obtain a MAPE of 3.8%, while the experimental data exhibit an averaged relative error margin of 1.2%Dréau et al. (2012). The obtained MAPE of the improved theoretical values for data set II. is 1.5%, which is better than the relative error margin of the experimental data of 3.3%Taminiau et al. (2012). This suggests that the theoretical values are ”overfitted” for data set II., i.e. multiple matching hyperfine values were found within the error margin of the experimental data. By selecting the closest ones, we could obtain a MAPE smaller than the experimental error bar. The most accurate hyperfine values, with a relative error margin of , together with nuclear spin positions are provided for data set III. in Ref van de Stolpe et al. (2023). These high-precision measurements allow us to make a reliable assessment of the error bar of the theoretical values. Considering the 29 most accurately measured and positioned nuclear spins, we obtain a MAPE of 1.79%. In Fig. (3), we also depicted the ARE of hyperfine values obtained with the PBE functional. As can be seen, the HSE(0.25) functional consistently improves on the PBE values, especially for nuclear spins found close to the NV center. We also note that the use of accurate DFT spin density has high relevance. Considering point spin density approximation, i.e. where is the center of the NV center, we obtain a mean absolute relative error (MARE) of 76% for data set III. see Supplementary Information.
@@
Finally, we investigate the source of residual errors and possible further measures to improve the computed hyperfine tensors. For this study, we use data set III. First of all, it is notable that the mean signed relative error (MSRE) of -0.12% is significantly smaller than the MARE of 1.79%, suggesting that there are no large systematic errors and the discrepancies may be related to numeral uncertainties. Next, we plot the MARE as a function of the Fermi contact term, see Fig. (4), and as a function of the distance, see Supplementary Information. There seems to be a correlation between the largest errors and the value of the Fermi contact term. Due to the distance dependence of the Fermi contact term, the MARE seems to decay with the distance of the nuclear spins, see Supplementary Information. It should be noted; however, that large errors are also obtained for a few nuclear spins that are beyond the boundaries of the supercell, where we explicitly neglect the Fermi contact terms. Here, the neglect of the Fermi contact term may be the source of the error. To attempt reducing the errors, we tune the mixing parameter of the HSE functional and study the variation of the MARE obtained for data set III., see Supplementary Information. By reducing the mixing parameter from 0.25 to 0.2, we obtain a slightly decreased MARE of 1.69%, although an enlarged MSRE of -0.32%. These results indicate that tuning the functional’s inner parameters may lead to an improved description of the spin density, which in turn can reduce the relative error between theory and experiments. Overall, the hyperfine values may be further improved by enhancing numerical accuracy for the core/Fermi contact contribution and by fine tuning the functional.
@@
In summary, we demonstrated in this Letter that high-accuracy, finite-size effect-free hyperfine tensors can be calculated using an improved integration method. Compared to the industry standard VASP code, we could achieve a 100 fold reduction of the MARE of the theoretical values. Having more experimental data of high precision and further improved numerical accuracy in larger supercells may help to obtain superior theoretical hyperfine data compared to what has been presented here. The obtained and potentially improved future data is available online under Ref. [33]. The provided data can be used for high-precision simulation of the NV center-nuclear spin few-body quantum systems as well as positioning nuclear spin around the NV center.
@@ Methods
For VASP calculations, we use a 512-atom and a 1728-atom supercell of diamond including a single NV center in the middle of the supercell. For the calculations, we use PBE and HSE06 exchange-correlation functional, 500 eV cutoff energy of the plane-wave basis set, -point sampling of the Brillouin zone, and high convergence criteria (PREC = Accurate). The energy threshold for the self-consistent field calculations is set to 10 eV. The structure of the defect is optimized as far as the largest force is smaller than 10 eV/Å. For the real-space hyperfine tensor calculations, we use the convergent spin density of the 1728-atom supercell model defined on a fine real-space grid of 0.036 Å (/600) spacing obtained by VASP.
@@ Acknowledgments
Insightful comments from Alexander P. Nizovtsev are highly appreciated. This work was supported by the National Research, Development and Innovation Office of Hungary (NKFIH) within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004) and within projects FK 135496 and FK 145395. V.I. acknowledges support from the Knut and Alice Wallenberg Foundation through the WBSQD2 project (Grant No. 2018.0071). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at the Swedish National Infrastructure for Computing (SNIC) at Tetralith, partially funded by the Swedish Research Council through grant agreement no. 2022-06725.

### v3

@@ Introduction
Hyperfine coupling tensors can be calculated using different first principles methods, such as density functional theory (DFT) and wave functions-based methods. The accuracy of the computed hyperfine tensors for close nuclear spins is remarkable. For example, considering paramagnetic point defects in semiconductors a mean absolute relative error of 4.7% has been reported.Szász et al. (2013) Since the hyperfine structure of the point defect’s spin sublevels is unique like a bar code, one can compare the measured and computed hyperfine tensors to identify paramagnetic defects in semiconductors unambiguously, see for instance Refs. Son et al. ; Isoya et al. ; Ivády et al. (2017).
@@
As demonstrated recently by a supercell-size-scaling test in Ref. Ivády et al. (2018), the accuracy of the computed hyperfine parameters is only limited for the closest nuclear spin in 1-5 Å distances from the defect. The relative error sharply increases for nuclear spins located at large distances. As discussed in Ref. Ivády et al. (2018), this is presumably due to the periodic boundary condition and related finite size effects. Using finite cluster models, such as the C291H172 cluster and the C510H252 cluster used in Refs. Nizovtsev et al. (2014, ) to calculate the hyperfine coupling tensors at 1.5-8 Å distances from the defect, errors related to the periodic boundary condition can be eliminated. Another limitation of first principles calculations is the model size that maximizes the number of lattice sites that can be considered.
@@
In this work, we first demonstrate the inaccuracy of the numerical hyperfine parameters obtained with the industry standard VASP codeKresse and Hafner (1994); Kresse and Furthmüller (1996). To resolve the underlying methodological issues we introduce a real-space integration method and the use of a large support lattice for considering nuclear spins outside the boundaries of the supercell. To benchmark our method, we carry out large-scale calculations for the NV center in diamond using different exchange-correlation functionals and compare the numerical results with available experimental hyperfine data sets. From the comparison, we conclude that the HSE06Heyd et al. (2003, 2006) functional with 0.2 mixing parameter performs best for the NV center in diamond resulting in a mean absolute percentage error of 1.7% for nuclear spins 6-30 Å distances from the NV center. This is a significant improvement compared to previous theoretical predictions obtained by using VASP. We show that the residual errors are likely related to the inaccurate calculation of the Fermi contact term. High accuracy hyperfine tensors of lattice sites as well as volumetric hyperfine data with Å spatial resolution are published together with this article 111Theoretical hyperfine tensors for all lattice sites within 30 Å distance from the NV center are available at https://ivadygroup.elte.hu/hyperfine and ready to be used for modeling NV center quantum nodes and positioning nuclear spins around the NV center in diamond in nano-NMR measurements.
@@ I Results and Discussion
where is the hyperfine tensor and and are the electron spin and the nuclear spin operator vectors with and quantum numbers for the NV center and adjacent 13C nuclear spins of the diamond lattice, respectively. The two dominant contributions to the hyperfine tensor are the Fermi contact interaction, , and the magnetic dipole-dipole coupling of the electron and nuclear spins, .
@@
First, we use this method as implemented in VASP to calculate hyperfine tensors of all sites in a 512-atom and a 1728-atom supercell of diamond including a single NV center in the middle of the supercell, see Fig. 1. For the calculations, we use HSE06 exchange-correlation functional, 500 eV cutoff energy of the plane-wave basis set, -point sampling of the Brillouin zone, and high convergence criteria. The structure of the defect is optimized as far as the largest force is smaller than 10-3 eV/Å. To compare our results with the experimental values, see Fig. 2(a), we either compute the hyperfine splitting as Dréau et al. (2012), where are the elements of the hyperfine tensor, which is compared with experimental values in data set I. taken from Ref. Dréau et al. (2012), or compare the hyperfine tensor element with experimental values obtained from high precision measurements in Ref. Taminiau et al. (2012) (data set II.) and Ref. van de Stolpe et al. (2024) (data set III.). Note that the latter data set is a refined and extended version of the data published in Ref. Abobeih et al. (2019). From the reported 50 nuclear spins, we consider only those for which both the position and the hyperfine parameter are reported with high accuracy, see Fig. 2(a). When comparing the computed values with data sets II.-III., we found a consistent sign difference between theory and experiment. The sign of our hyperfine values agrees with other theoretical calculations, i.e. VASP’s values as well as the values reported in Refs. Nizovtsev et al. (2014, ). Therefore, we anticipate that the discrepancy originates from the convention used in the experiments. Hereinafter, we omit this sign difference.
@@
The hyperfine splitting values computed with our method are significantly improved compared to the currently available implementationSzász et al. (2013). The relative error of the theoretical values reduce from to , see Fig. 2(b) and Fig. 3. For different data sets, we obtain different mean absolute percentage errors (MAPE). For data set I. we obtain a MAPE of 3.8%, while the experimental data exhibit an averaged relative error margin of 1.2%Dréau et al. (2012). The obtained MAPE of the improved theoretical values for data set II. is 1.5%, which is better than the relative error margin of the experimental data of 3.3%Taminiau et al. (2012). This suggests that the theoretical values are ”overfitted” for data set II., i.e. multiple matching hyperfine values were found within the error margin of the experimental data. By selecting the closest ones, we could obtain a MAPE smaller than the experimental error bar. The most accurate hyperfine values, with a relative error margin of , together with nuclear spin positions are provided for data set III.van de Stolpe et al. (2024). These high-precision measurements allow us to make a reliable assessment of the error bar of the theoretical values. Considering the 29 most accurately measured and positioned nuclear spins, we obtain a MAPE of 1.79%. In Fig. (3), we also depicted the ARE of hyperfine values obtained with the PBE functional. As can be seen, the HSE06 functional consistently improves on the PBE values, especially for nuclear spins found close to the NV center. We also note that the use of accurate DFT spin density has high relevance. Considering point spin density approximation, i.e. where is the center of the NV center, we obtain a mean absolute relative error (MARE) of 76% for data set III. see Supplementary Figure S1.
@@
Finally, we investigate the source of residual errors and possible further measures to improve the computed hyperfine tensors. For this study, we use data set III. First of all, it is notable that the mean signed relative error (MSRE) of -0.12% is significantly smaller than the MARE of 1.79%, suggesting that there are no large systematic errors and the discrepancies may be related to numeral uncertainties. Next, we plot the MARE as a function of the Fermi contact term, see Fig. (4), and as a function of the distance, see Supplementary Figure S2. There seems to be a correlation between the largest errors and the value of the Fermi contact term. Due to the distance dependence of the Fermi contact term, the MARE seems to decay with the distance of the nuclear spins, see Supplementary Figure S2. It should be noted; however, that large errors are also obtained for a few nuclear spins that are beyond the boundaries of the supercell, where we explicitly neglect the Fermi contact terms. Here, the neglect of the Fermi contact term may be the source of the error. To attempt reducing the errors, we tune the mixing parameter of the HSE functional and study the variation of the MARE obtained for data set III., see Supplementary Figure S3. By reducing the mixing parameter from 0.25 to 0.2, we obtain a slightly decreased MARE of 1.69%, although an enlarged MSRE of -0.32%. These results indicate that tuning the functional’s inner parameters may lead to an improved description of the spin density, which in turn can reduce the relative error between theory and experiments. Overall, the hyperfine values may be further improved by enhancing numerical accuracy for the core/Fermi contact contribution and by fine-tuning the functional.
@@
In summary, we demonstrated in this article that high-accuracy, finite-size effect-free hyperfine tensors can be calculated using an improved integration method. Compared to the industry standard VASP code, we could achieve a 100 fold reduction of the MARE of the theoretical values. Having more experimental data of high precision and further improved numerical accuracy in larger supercells may help to obtain superior theoretical hyperfine data compared to what has been presented here. The obtained and potentially improved future data is available online [34]. The provided data can be used for high-precision simulation of the NV center-nuclear spin few-body quantum systems as well as positioning nuclear spin around the NV center.
@@ Methods
For VASP calculations, we use a 512-atom and a 1728-atom supercell of diamond including a single NV center in the middle of the supercell. For the calculations, we use the Perdew–Burke–Ernzerhof (PBE)Perdew et al. (1996) and the Heyd–Scuseria–Ernzerhof (HSE06)Heyd et al. (2003, 2006) exchange-correlation functional, 500 eV cutoff energy of the plane-wave basis set, -point sampling of the Brillouin zone, and high convergence criteria (PREC = Accurate). The energy threshold for the self-consistent field calculations is set to 10-6 eV. The structure of the defect is optimized as far as the largest force is smaller than 10-3 eV/Å. For the real-space hyperfine tensor calculations, we use the convergent spin density of the 1728-atom supercell model defined on a fine real-space grid of 0.036 Å (/600) spacing obtained by VASP.
@@ Acknowledgments
Insightful comments from Alexander P. Nizovtsev are highly appreciated. This work was supported by the National Research, Development and Innovation Office of Hungary (NKFIH) within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004) and within project FK 145395. V.I. acknowledges support from the Knut and Alice Wallenberg Foundation through the WBSQD2 project (Grant No. 2018.0071). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at the Swedish National Infrastructure for Computing (SNIC) at Tetralith, partially funded by the Swedish Research Council through grant agreement no. 2022-06725.